OpenStudy (anonymous):

5+(square root)3/ 2- (square root)3

OpenStudy (anonymous):

what is the actual problem...? \[5+\sqrt{3}/2 - \sqrt{3}=... or (5+\sqrt{3})/2 -\sqrt{3}...?\]

OpenStudy (anonymous):

the second one